pablos coffee shop makes a blend of coffee that is a mixture of two types of coffee. type A coffee cpst pablo $5.50 per pound, and type B coffee costs $4.30 per pound. this month, pavlo made 127 pounds of the blend, for a total cost of $607.30. how many pounds of type A coffee did he use?
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

A+B= 127 subtract B from both sides:
A= 127-B We'll use this value for A next.
5.50A+4.30B = 607.30 Now substitute for A:
5.50(127-B)+4.30B = 607.30
698.5-5.50B+4.30B = 607.30 Now subtract 698.5 from both sides and subtract B on left:
-1.2B = -91.20 Now divide both sides by -1.2 (and remember: -/- = +)
B = 76 pounds. And:
127-76 = 51 pounds are type A

Answer by ikleyn(52847)

.
pablos coffee shop makes a blend of coffee that is a mixture of two types of coffee. type A coffee cpst pablo $5.50 per pound, and type B coffee costs $4.30 per pound. this month, pavlo made 127 pounds of the blend, for a total cost of $607.30. how many pounds of type A coffee did he use?
------------------------------------------------------------------

See the lesson Word problems on mixtures for dry substances like coffee beans, nuts, cashew and peanuts in this site.


Very similar problem (Problem 2) was solved there for you.

Consider it as a sample. Read it attentively, and then solve your problem by substituting your data.